OpenStudy (anonymous):

How many moles of copper are in 6,000,000 atoms of copper?

OpenStudy (anonymous):

just take 6000000 and multiply it byt the atomic mass

OpenStudy (anonymous):

1 copper has 63.54 mol/g

OpenStudy (abb0t):

Wrong @Maisa101 you are using dimensional analysis to cancel out the number of atoms given by avogradro's number: \(6.022 \times 10^{23} \large \frac{mole}{atom}\) That means you are DIVIDING, not multiplying.

OpenStudy (anonymous):

hmmm, i'm sorry i'm a little confused what's dimensional analysis?

OpenStudy (abb0t):

Same as what you did with fractions in algebra, except the process is called "dimensional analysis" in chemistry. Crossing out like terms to get the desired product.

OpenStudy (anonymous):

ahh, I see then. Thanks for correcting me, I would hate to give someone the wrong answer.

OpenStudy (whpalmer4):

1 mole of copper atoms contains Avogadro's number (\(N_A =6.02214179*10^{23}\)) atoms. Clearly, 6,000,000 atoms is only going to give us a small fraction of a mole!

OpenStudy (abb0t):

It might not be as clear to them as it is for you. , otherwise they wouldn't be here seeking help :)

OpenStudy (anonymous):

Thank you all for the help :) I did get it eventually :D